# hyperExp head to column names, separator = comma, centered tabular = —c—c—c—c—c—c—c—c—c—c—, table head =     FW   FW EleSym   AGM   AGM EleSym Error Rel. time S(%) Rel. time S(%) Rel. time S(%) Rel. time S(%) , late after line =

### v3

@@ 1 Introduction
Hyperbolicity cones [gaarding1959inequality, guler1997hyperbolic, BGLS01, renegar2004hyperbolic] are a far-reaching family of closed convex cones containing all symmetric cones and all polyhedral cones. In particular, the cone of $n\times n$ real symmetric positive semidefinite matrices ${\mathcal{S}^{n}_{+}}$ is a hyperbolicity cone. One distinctive feature of ${\mathcal{S}^{n}_{+}}$ is that the orthogonal projection onto ${\mathcal{S}^{n}_{+}}$ has a well-known expression that can be described in terms of the spectral decomposition of a matrix. More precisely, if $X$ is an $n\times n$ symmetric matrix, $v_{1},\ldots v_{n}$ are $n$ orthonormal eigenvectors of $X$ and $\lambda_{1},\ldots,\lambda_{n}$ are the corresponding eigenvalues, then the projection onto ${\mathcal{S}^{n}_{+}}$ with respect to the Frobenius norm is the result of “zeroing the negative eigenvalues in the spectral decomposition”:
$P_{{\mathcal{S}^{n}_{+}}}(X)=\operatorname*{arg\,min}_{Y\in{\mathcal{S}^{n}_{+}}}\left\|Y-X\right\|_{F}=\sum_{i=1}^{n}\max(0,\lambda_{i})v_{i}v_{i}^{\top},$ (1.1)
where $\left\|\cdot\right\|_{F}$ is the Frobenius norm. Analogously, for a hyperbolicity cone there is a natural notion of eigenvalues (see Section 2.1) that is strong enough to allow the extension of certain linear algebraic results about symmetric matrices, e.g., [BGLS01, renegar2004hyperbolic]. With this in mind, the starting point of this project was the following questions:
Given an arbitrary hyperbolicity cone $\Lambda\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, how to compute the projection of $x$ onto $\Lambda$ efficiently, i.e., how to compute $P_{\Lambda}(x)\coloneqq\operatorname*{arg\,min}_{y\in\Lambda}\left\|y-x\right\|$? Are there closed form expressions analogous to (1.1) in term of the eigenvalues of $x$
First, we mention briefly why these questions are important. Besides being an interesting geometric question on its own, the usefulness of having a readily available projection operator for some convex set is well-documented in optimization (e.g., see [HM11]). It is, after all, a basic requirement for the applicability of several algorithms. For example, augmented Lagrangian methods for solving conic optimization problems typically require the projection operator onto the underlying cone or its dual, e.g., see [SS04, Section 2]. The same is true for operator splitting approaches, e.g., see [OCPB16, Section 3] and [BNZ24, Section 3.2] which discuss specifically the case of conic constraints.
@@ 2 Preliminaries
We write +$\mathbb{R}^{n}_{+}$ for the nonnegative orthant, i.e., the elements $u\in\mathbb{R}^{n}$ such that $u_{i}\geq 0$ for every $i$.
[…]
@@
[…]
@@
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|,\qquad\forall x,y\in\mathbb{R}^{n}.$
@@
$f(\theta x+(1-\theta)y)\leq\theta f(x)+(1-\theta)f(y)-\frac{\mu}{2}\theta(1-\theta)\left\|x-y\right\|^{2},\qquad\forall x,y\in\mathbb{R}^{n}.$
@@ 2.1 Hyperbolic polynomials
Suppose that $p$ has degree $d$. We define the map $\lambda:\mathbb{R}^{n}\to\mathbb{R}^{d}_{\downarrow}$ that maps $x\in\mathbb{R}^{n}$ to the $d$ roots of the polynomial $p(x-te)$, ordered from largest to smallest. That is, we have
@@
see Section 2 in [renegar2004hyperbolic]. If $p$ and $e$ are clear from the context, we write $\Lambda(p,e)$ as $\Lambda$. If $\hat{e}\in\mathrm{int}\,\Lambda(p,e)$, then $\Lambda(p,e)=\Lambda(p,\hat{e})$ holds, see [renegar2004hyperbolic, Theorem 3].
@@
$D_{e}p(x)=\lim_{t\to 0}\frac{p(x+te)-p(x)}{t},\quad\forall x\in\mathbb{R}^{n}.$
@@
$\partial^{r}\Lambda^{(1)}=\partial^{r+1}\Lambda$
@@
$\partial^{1}\Lambda^{(1)}\cap\Lambda=\partial^{2}\Lambda$
@@ 2.2 Generalized minimum eigenvalue functions
First we observe that $\lambda_{\min}(x)=\inf\,\{t\mid x-te\not\in{\mathcal{K}}\}$ holds and this was the original definition considered in [Renegar16, Section 2], see also [L17, Section 2.5] and [IL17, Section 2]. Given that there is a dependency on $e$ and ${\mathcal{K}}$, it might be appropriate to use some notation similar to “${\lambda^{{\mathcal{K}},e}_{\min}}$”, but since there will be no ambiguity regarding the chosen ${\mathcal{K}}$ and $e$, we will use the simpler $\lambda_{\min}$. We recall the following basic properties of $\lambda_{\min}$.
@@ 3. $(iii)$
Since $e$ is an interior point of ${\mathcal{K}}$, there exists $u>0$ such that $e+ux\in{\mathcal{K}}$ and since ${\mathcal{K}}$ is a cone, $x+e/u\in{\mathcal{K}}$ holds as well. This shows that $\lambda_{\min}(x)>-\infty$. On the other hand if $\lambda_{\min}(x)=\infty$, then $x-te$ belongs to ${\mathcal{K}}$ for all $t$. This would imply that $-e\in{\mathcal{K}}$ (e.g., see [rockafellar, Theorem 8.3]), which would contradict the pointedness of ${\mathcal{K}}$. This shows that item $(i)$ holds.
@@
$\lambda(x-te)=\lambda(x)-t\textbf{1},\qquad\forall t\in\mathbb{R},$
where $\textbf{1}\in\mathbb{R}^{d}$ is the vector where each component is $1$. Recalling that $x-te\in\Lambda(p,e)$ if and only if $\lambda(x-te)\geq 0$, we see that the condition “$x-te\in\Lambda(p,e)$” implies the component-wise inequality
@@
In particular, the maximum value that $t$ can assume under the constraint “$x-te\in\Lambda(p,e)$” is $\lambda_{d}(x)$, where $\lambda_{d}(x)$ is the smallest root of $t\mapsto p(x-te)$. Conversely, if $t\coloneqq\lambda_{d}(x)$, we have $\lambda_{d}(x-\lambda_{d}(x)e)=0$, so $x-\lambda_{d}(x)e\in\Lambda(p,e)$. That is, $\lambda_{d}(x)$ is the optimal solution to the maximization problem in (2.4). ∎
@@ 2.3 The Frank-Wolfe Method
Let ${x}_{\mathrm{opt}}$ denote an optimal solution to (2.5). Since $f$ is convex, for $x\in C$ we have $\langle-\nabla f(x),{x}_{\mathrm{opt}}-x\rangle\geq f(x)-f({x}_{\mathrm{opt}})\geq 0$. We also recall that $x\in C$ is optimal if and only if $-\nabla f(x)$ belongs to the normal cone of $C$ at $x$. Consequently, the FW gap has the following properties for $x\in C$, see also [bomze2021frank, Section 5.1]:
  • $G(x)$ is always nonnegative and equal to 00 if and only if $x$ is optimal.
@@ 3 Distance function and projection under the norm induced by a hyperbolic polynomial
The nonnegative orthant +$\mathbb{R}^{n}_{+}$ can be realized as a hyperbolicity cone by taking $p:\mathbb{R}^{n}\to\mathbb{R}$ to be the polynomial $p(x)\coloneqq\prod_{i=1}^{n}x_{i}$ and $e\coloneqq(1,\ldots,1)$. With that, the induced inner product and norm are the usual Euclidean ones. A proof that $p$ is isometric is given in [BGLS98_pre, Section 7].
@@
We present a short argument that $\det$ is isometric, see also [BGLS98_pre, Section 7]. Let $y,z\in\mathcal{S}^{n}$ and consider the spectral decomposition of $y$ so that $y=u{\mathrm{diag}\,}(\lambda(y))u^{\top}$ holds, where $u$ is an $n\times n$ orthogonal matrix, ${\mathrm{diag}\,}(\lambda(y))$ is the diagonal matrix associated to $\lambda(y)$ and $u^{\top}$ is the transpose of $u$. Then, letting $x\coloneqq u{\mathrm{diag}\,}(\lambda(z))u^{\top}$, we have $\lambda(x)=\lambda(z)$ and $x+y=u{\mathrm{diag}\,}(\lambda(x)+\lambda(y))u^{\top}$ so that $\lambda(x+y)=\lambda(x)+\lambda(y)$ holds.
@@ 3.1 The distance function to a hyperbolicity cone
Let $x\in\mathbb{R}^{n}$. Because $x$ belongs to $\Lambda$ if and only if $\lambda(x)\in\mathbb{R}^{d}_{+}$, a reasonable guess is that ${\mathrm{dist}\,}(x,\Lambda)$ satisfies
@@
where $\lambda(x)^{+}$ is the projection of $\lambda(x)$ onto +$\mathbb{R}^{d}_{+}$, which is obtained by zeroing the negative components of $\lambda(x)$. Since $\lambda(x)\in\mathbb{R}^{d}_{\downarrow}$ holds, we also have $\lambda(x)^{+}\in\mathbb{R}^{d}_{\downarrow}$. We conclude that
@@
We now take a look at the requirement that $\lambda(\Lambda)$ coincides with +$\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$. By construction, $\lambda(\Lambda)$ is always contained in +$\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$, so the nontrivial part is the opposite containment. As we will see in the next lemma, the condition $\lambda(\Lambda)=\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$ is equivalent to $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ when $p$ is isometric. The significance of the equality $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ is that it makes it possible to prove certain results on the Fenchel conjugate of spectral functions, see [BGLS01, Theorem 5.5]. The equality $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ also indicates that every element in $\mathbb{R}^{d}_{\downarrow}$ is an eigenvalue vector of some $x\in\mathbb{R}^{n}$.
@@
Since $u^{+}\in\mathbb{R}^{d}_{+}$ as well, by hypothesis, there exists $y\in\Lambda$ such that $\lambda(y)=u^{+}$. Similarly, there is $z\in\Lambda$ such that $\lambda(z)=(-u^{-})^{\downarrow}$. Since the eigenvalues of $-z$ are the negatives of the eigenvalues of $z$, we have $\lambda(-z)=u^{-}$.
@@ 3.2 Partial results on the projection operator
By definition, $f$ is differentiable and $\nabla f$ is the projection operator onto +$\mathbb{R}^{n}_{+}$, i.e.
@@ 2. $(ii)$
The determinant polynomial over the $n\times n$ symmetric matrices $\det:\mathcal{S}^{n}\to\mathbb{R}$ is diagonalizable in the sense of Definition 3.8 and a suitable $\Delta$ is given by the map that takes a vector in $\mathbb{R}^{n}$ and maps to the corresponding diagonal matrix in $\mathcal{S}^{n}$. In this case, $\mathcal{G}_{\det}$ corresponds to the linear transformations on $\mathcal{S}^{n}$ that preserve the eigenvalue map. It is a non-trivial but well-known fact that every such transformation must be of the form $x\mapsto vxv^{\top}$ for some $n\times n$ orthogonal matrix $v$, where $v^{\top}$ indicates the usual matrix transpose. This follows, for example, from the results contained in [Li79, Section 5]. Overall, for the case of the determinant polynomial $\det:\mathcal{S}^{n}\to\mathbb{R}$, Definition 3.8 expresses the well-known fact that every symmetric matrix has a spectral decomposition and the linear function that takes a vector in $\mathbb{R}^{n}$ to its corresponding diagonal matrix in $\mathcal{S}^{n}$ is an isometry. For other examples of diagonalizable hyperbolic polynomials see [BGLS98_pre, Section 7]. Diagonalizable hyperbolic polynomials must be isometric by [BGLS98_pre, Theorem 6.5].
@@
where $\lambda(x)_{+}$ is the projection of $\lambda(x)$ onto +$\mathbb{R}^{n}_{+}$.

### v4

# arXiv:2407.09213v4[math.OC] 03 Mar 2026

@@ 1 Introduction
Hyperbolicity cones [gaarding1959inequality, guler1997hyperbolic, BGLS01, renegar2004hyperbolic] are a far-reaching family of closed convex cones containing all symmetric cones and all polyhedral cones. In particular, the cone of $n\times n$ real symmetric positive semidefinite matrices $\mathcal{S}^{n}_{+}$} is a hyperbolicity cone. One distinctive feature of $\mathcal{S}^{n}_{+}$} is that the orthogonal projection onto $\mathcal{S}^{n}_{+}$} has a well-known expression that can be described in terms of the spectral decomposition of a matrix. Analogously, for a hyperbolicity cone there is a natural notion of eigenvalues (see Section 2.1) that is strong enough to allow the extension of certain linear algebraic results about symmetric matrices, e.g., [BGLS01, renegar2004hyperbolic]. With this in mind, the starting point of this project was the following questions:
Given an arbitrary hyperbolicity cone $\Lambda\subseteq\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, how to compute the projection of $x$ onto $\Lambda$ efficiently, i.e., how to compute $P_{\Lambda}(x)\coloneqq\operatorname*{arg\,min}_{y\in\Lambda}\left\|y-x\right\|$? Are there closed form expressions in terms of the eigenvalues of $x$ ?
In a nutshell our answers are as follows. We only found a closed form expression for the projection operator for a rather narrow family of hyperbolicity cones (see Section 3 and Proposition 3.9). For more general cones, numerical methods seem necessary, which we discuss in detail in this paper in Sections 4 and 5.
Before we move on, we take a step back to motivate these questions. Besides being an interesting geometric question on its own, the usefulness of having a readily available projection operator for some convex set is well-documented in optimization (e.g., see [HM11]). It is, after all, a basic requirement for the applicability of several algorithms. For example, augmented Lagrangian methods for solving conic optimization problems typically require the projection operator onto the underlying cone or its dual, e.g., see [SS04, Section 2]. The same is true for operator splitting approaches, e.g., see [OCPB16, Section 3] and [BNZ24, Section 3.2] which discuss specifically the case of conic constraints.
@@ 2 Preliminaries
We write $\mathbb{R}^{n}_{+}$ for the nonnegative orthant, i.e., the elements $u\in\mathbb{R}^{n}$ such that $u_{i}\geq 0$ for every $i$.
For a convex subset $S\subseteq\mathbb{R}^{n}$, we denote its indicator function, recession cone, interior and relative interior by $\delta_{S}$, $0^{+}S$, $\mathrm{int}\,S$ and $\mathrm{ri}\,S$, respectively. Additionally, we suppose $\mathbb{R}^{n}$ is endowed with an inner product $\langle\cdot,\cdot\rangle$ and a corresponding induced norm $\left\|\cdot\right\|$. With that, we denote by $S^{\perp}$ the set of elements orthogonal to $S$. For a convex cone ${\mathcal{K}}\subseteq\mathbb{R}^{n}$, we define its dual cone as
@@
where by a slight abuse of notation we use the same symbol $\|\cdot\|$ to indicate the underlying norm in $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$.
@@
$\left\|\nabla f(x)-\nabla f(y)\right\|\leq L\left\|x-y\right\|,\hskip 20.00003pt\forall x,y\in\mathbb{R}^{n}.$
@@
$f(\theta x+(1-\theta)y)\leq\theta f(x)+(1-\theta)f(y)-\frac{\mu}{2}\theta(1-\theta)\left\|x-y\right\|^{2},\hskip 20.00003pt\forall x,y\in\mathbb{R}^{n}.$
@@ 2.1 Hyperbolic polynomials
Suppose that $p$ has degree $d$. We define the map $\lambda:\mathbb{R}^{n}\to\mathbb{R}^{d}_{\downarrow}$ that maps $x\in\mathbb{R}^{n}$ to the $d$ roots of the polynomial p(xte)p(x-te), ordered from largest to smallest. That is, we have
@@
[…]
@@
$D_{e}p(x)=\lim_{t\to 0}\frac{p(x+te)-p(x)}{t},\qquad\forall x\in\mathbb{R}^{n}.$
@@
$\partial^{r}\Lambda^{(1)}=\partial^{r+1}\Lambda.$
@@
$\partial^{1}\Lambda^{(1)}\cap\Lambda=\partial^{2}\Lambda.$
@@ 2.2 Generalized minimum eigenvalue functions
First we observe that $\lambda_{\min}(x)=\inf\,\{t\mid x-te\not\in{\mathcal{K}}\}$ holds and this was the original definition considered in [Renegar16, Section 2], see also [L17, Section 2.5] and [IL17, Section 2]. Given that there is a dependency on $e$ and $\mathcal{K}$}, it might be appropriate to use some notation similar to $\lambda^{{\mathcal{K}},e}_{\min}$}”, but since there will be no ambiguity regarding the chosen $\mathcal{K}$} and $e$, we will use the simpler $\lambda_{\min}$. We recall the following basic properties of $\lambda_{\min}$.
@@ 3. $(iii)$
Since $e$ is an interior point of $\mathcal{K}$}, there exists $u>0$ such that $e+ux\in{\mathcal{K}}$ and since $\mathcal{K}$} is a cone, $x+e/u\in{\mathcal{K}}$ holds as well. This shows that $\lambda_{\min}(x)>$-$\infty$. On the other hand if $\lambda_{\min}(x)=\infty$, then xtex-te belongs to $\mathcal{K}$} for all $t$. This would imply that $e\in{\mathcal{K}}$ (e.g., see [rockafellar, Theorem 8.3]), which would contradict the pointedness of $\mathcal{K}$}. This shows that item $(i)$ holds.
@@
$\lambda(x-te)=\lambda(x)-t\textbf{1},\hskip 20.00003pt\forall t\in\mathbb{R},$
where $\textbf{1}\in\mathbb{R}^{d}$ is the vector where each component is $1$. Recalling that $x-te\in\Lambda(p,e)$ if and only if $\lambda(x-te)\geq 0$, we see that the condition $x-te\in\Lambda(p,e)$” implies the component-wise inequality
@@
In particular, the maximum value that $t$ can assume under the constraint $x-te\in\Lambda(p,e)$” is $\lambda_{d}(x)$, where $\lambda_{d}(x)$ is the smallest root of $t\mapsto p(x-te)$. Conversely, if $t\coloneqq\lambda_{d}(x)$, we have $\lambda_{d}(x-\lambda_{d}(x)e)=0$, so $x-\lambda_{d}(x)e\in\Lambda(p,e)$. That is, $\lambda_{d}(x)$ is the optimal solution to the maximization problem in (2.4). ∎
@@ 2.3 The Frank-Wolfe Method
Let ${x}_{\mathrm{opt}}$ denote an optimal solution to (2.5). Since $f$ is convex, for $x\in C$ we have $\langle-\nabla f(x),{x}_{\mathrm{opt}}-x\rangle\geq f(x)-f({x}_{\mathrm{opt}})\geq 0$. We also recall that $x\in C$ is optimal if and only if $\nabla f(x)$ belongs to the normal cone of $C$ at $x$. Consequently, the FW gap has the following properties for $x\in C$, see also [bomze2021frank, Section 5.1]:
  • $G(x)$ is always nonnegative and equal to 0 if and only if $x$ is optimal.
@@ 3 Distance function and projection under the norm induced by a hyperbolic polynomial
The nonnegative orthant $\mathbb{R}^{n}_{+}$ can be realized as a hyperbolicity cone by taking $p:\mathbb{R}^{n}\to\mathbb{R}$ to be the polynomial $p(x)\coloneqq\prod_{i=1}^{n}x_{i}$ and $e\coloneqq(1,\ldots,1)$. With that, the induced inner product and norm are the usual Euclidean ones. A proof that $p$ is isometric is given in [BGLS98_pre, Section 7].
@@
We present a short argument that $\det$ is isometric, see also [BGLS98_pre, Section 7]. Let $y,z\in\mathcal{S}^{n}$ and consider the spectral decomposition of $y$ so that $y=u{\mathrm{diag}\,}(\lambda(y))u^{\top}\hskip-1.0pt$ holds, where $u$ is an $n\times n$ orthogonal matrix, ${\mathrm{diag}\,}(\lambda(y))$ is the diagonal matrix associated to $\lambda(y)$ and $u^{\top}\hskip-1.0pt$ is the transpose of $u$. Then, letting $x\coloneqq u{\mathrm{diag}\,}(\lambda(z))u^{\top}\hskip-1.0pt$, we have $\lambda(x)=\lambda(z)$ and $x+y=u{\mathrm{diag}\,}(\lambda(x)+\lambda(y))u^{\top}\hskip-1.0pt$ so that $\lambda(x+y)=\lambda(x)+\lambda(y)$ holds.
@@
Before we move on, we recall that if $X$ is an $n\times n$ symmetric matrix, $v_{1},\ldots v_{n}$ are $n$ orthonormal eigenvectors of $X$ and $\lambda_{1},\ldots,\lambda_{n}$ are the corresponding eigenvalues, then the projection onto $\mathcal{S}^{n}_{+}$} with respect to the Frobenius norm is the result of “zeroing the negative eigenvalues in the spectral decomposition”:
$P_{{\mathcal{S}^{n}_{+}}}(X)=\operatorname*{arg\,min}_{Y\in{\mathcal{S}^{n}_{+}}}\left\|Y-X\right\|_{F}=\sum_{i=1}^{n}\max(0,\lambda_{i})v_{i}v_{i}^{\top}\hskip-1.0pt,$ (3.2)
where $\left\|\cdot\right\|_{F}$ is the Frobenius norm. This implies that the distance function with respect the Frobenius norm is given by
${\mathrm{dist}\,}(x,{\mathcal{S}^{n}_{+}})^{2}=\sum_{i=1}^{n}\min(\lambda_{i},0)^{2}.$ (3.3)
In the remainder of this section, by imposing increasingly more restrictive assumptions, we will discuss potential analogues of (3.2) and (3.3) for the case of the hyperbolicity cones.
@@ 3.1 The distance function to a hyperbolicity cone
Let $x\in\mathbb{R}^{n}$. Because $x$ belongs to $\Lambda$ if and only if $\lambda(x)\in\mathbb{R}^{d}_{+}$, in analogy to (3.3), a reasonable guess is that ${\mathrm{dist}\,}(x,\Lambda)$ satisfies
@@
where $\lambda(x)^{+}$ is the projection of $\lambda(x)$ onto $\mathbb{R}^{d}_{+}$, which is obtained by zeroing the negative components of $\lambda(x)$. Since $\lambda(x)\in\mathbb{R}^{d}_{\downarrow}$ holds, we also have $\lambda(x)^{+}\in\mathbb{R}^{d}_{\downarrow}$. We conclude that
@@
We now take a look at the requirement that $\lambda(\Lambda)$ coincides with $\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$. By construction, $\lambda(\Lambda)$ is always contained in $\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$, so the nontrivial part is the opposite containment. As we will see in the next lemma, the condition $\lambda(\Lambda)=\mathbb{R}^{d}_{+}\cap\mathbb{R}^{d}_{\downarrow}$ is equivalent to $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ when $p$ is isometric. The significance of the equality $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ is that it makes it possible to prove certain results on the Fenchel conjugate of spectral functions, see [BGLS01, Theorem 5.5]. The equality $\lambda(\mathbb{R}^{n})=\mathbb{R}^{d}_{\downarrow}$ also indicates that every element in $\mathbb{R}^{d}_{\downarrow}$ is an eigenvalue vector of some $x\in\mathbb{R}^{n}$.
@@
Since $u^{+}\in\mathbb{R}^{d}_{+}$ as well, by hypothesis, there exists $y\in\Lambda$ such that $\lambda(y)=u^{+}$. Similarly, there is $z\in\Lambda$ such that $\lambda(z)=(-u^{-})^{\downarrow}$. Since the eigenvalues of z-z are the negatives of the eigenvalues of $z$, we have $\lambda(-z)=u^{-}$.
@@ 3.2 Partial results on the projection operator
By definition, $f$ is differentiable and $\nabla f$ is the projection operator onto $\mathbb{R}^{n}_{+}$, i.e.
@@ 2. $(ii)$
The determinant polynomial over the $n\times n$ symmetric matrices $\det:\mathcal{S}^{n}\to\mathbb{R}$ is diagonalizable in the sense of Definition 3.8 and a suitable $\Delta$ is given by the map that takes a vector in $\mathbb{R}^{n}$ and maps to the corresponding diagonal matrix in $\mathcal{S}^{n}$. In this case, $\mathcal{G}_{\det}$ corresponds to the linear transformations on $\mathcal{S}^{n}$ that preserve the eigenvalue map. It is a non-trivial but well-known fact that every such transformation must be of the form $x\mapsto vxv^{\top}\hskip-1.0pt$ for some $n\times n$ orthogonal matrix $v$, where $v^{\top}\hskip-1.0pt$ indicates the usual matrix transpose. This follows, for example, from the results contained in [Li79, Section 5]. Overall, for the case of the determinant polynomial $\det:\mathcal{S}^{n}\to\mathbb{R}$, Definition 3.8 expresses the well-known fact that every symmetric matrix has a spectral decomposition and the linear function that takes a vector in $\mathbb{R}^{n}$ to its corresponding diagonal matrix in $\mathcal{S}^{n}$ is an isometry. For other examples of diagonalizable hyperbolic polynomials see [BGLS98_pre, Section 7]. Diagonalizable hyperbolic polynomials must be isometric by [BGLS98_pre, Theorem 6.5].
@@
where $\lambda(x)_{+}$ is the projection of $\lambda(x)$ onto $\mathbb{R}^{n}_{+}$.